### v1

# NoTeeline: Supporting Real-Time Notetaking from Keypoints with Large Language Models

Abstract.
Video has become a popular media form for information sharing and consumption. However, taking notes while watching a video requires significant time and effort. To address this, we propose a novel interactive system, NoTeeline, for taking real-time, personalized notes. NoTeeline lets users quickly jot down key points (micronotes), which are automatically expanded into full-fledged notes that capture the content of the user’s micronotes and are consistent with the user’s writing style. In a within-subjects study (N=12), we found that NoTeeline helps users create high-quality notes that capture the essence of their micronotes with a higher factual correctness (93.2%) while accurately reflecting their writing style. While using NoTeeline, participants experienced significantly reduced mental effort, captured satisfactory notes while writing 47% less text, and completed notetaking with 43.9% less time compared to a manual notetaking baseline.
Large Language Model, Natural Language Interface, Personalization, Writing Assistant
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXccs: Human-centered computing User interface toolkitsccs: Human-centered computing Natural language interfaces
[…]
MyFrameroundcorner=3pt, linecolor=gray, outerlinewidth=0.5pt, innertopmargin=5pt, innerbottommargin=5pt, innerrightmargin=5pt, innerleftmargin=5pt, backgroundcolor=gray!8!white, frametitleaboveskip=
Figure 1. NoTeeline allows users to jot down key points (micronotes), which are automatically expanded into full notes consistent with the user’s writing style. The input and output are marked in purple and green, respectively. A user writes a (a) micronote while watching a video. NoTeeline captures the timestamp, retrieves the corresponding (b) video transcript, and expands the micronote into (1) full notes. During the expansion, NoTeeline uses the user’s (c) previous notes to infer their writing style. NoTeeline also generates (2) personalized cue questions, (3) summary, and (4) themes for note categorization.
[…]
1. Introduction
”The more content you try to capture during a lecture or a meeting, the less you’re thinking about what’s being said. You burn through most of your attention parroting the source.” - Ryder Carroll (Carroll, 2018)
Online video platform has emerged as a popular form of communication and information sharing (Tversky et al., 2002; Bétrancourt and Benetos, 2018). While accessing video content is easier than ever, the difficulty lies in capturing notes and organizing key information from these videos (Tabbers et al., 2004; Boyle, 2007). Effective note-taking is crucial for retaining and processing information, yet many users struggle with this task due to dual demands of absorbing new content and creating comprehensive notes simultaneously (Piolat et al., 2005; Jansen et al., 2017; Piolat, 2007). This often results in incomplete or poor-quality notes, or causes users to abandon note-taking altogether (Kafipour and Jafari, 2021; Flanigan and Titsworth, 2020; McCoy, 2020; Bui et al., 2013). By empowering users to take notes more easily, we can help them better engage with the video content.
Current note-taking tools often fall short in addressing the ongoing challenges users face, as they either require manual effort or generate generic, automated notes. Active note-taking tools have explored various methods to assist users in writing and organizing manual notes more efficiently (Subramonyam et al., 2020; Chiu et al., 1999; Cao et al., 2022). Despite offering features that speed up the process, these tools still necessitate manual input for creating comprehensive notes, which remains a significant challenge (Piolat et al., 2005). On the other hand, recent advancements in language processing and Large Language Models (LLMs) have led to tools that automatically generate notes or summaries from videos and live lectures (NoteGPT.io, 2024; Notion, 2024; Otter.ai, 2023). However, these automated tools often lack personalization and do not cater to individual user preferences (Biermann et al., 2022). This creates a gap between the manual, labor-intensive process of traditional note-taking and the impersonal nature of automatic note generation.
In this work, we aim to bridge these gaps by designing NoTeeline, a system that applies the principle of micronotes (Lin et al., 2004) to optimize the note-taking process. NoTeeline enables users to take less demanding micronotes with minimal effort and automatically expand them into full-fledged notes. By Micronotes, we refer to scribbles composed of couple-of-words and informal notations people typically use in quickly writing things down (Lin et al., 2004; Majewski, 2023). These micronotes are concise, causing minimal viewing distraction while capturing essential information densely. We developed a proof-of-concept interface, NoTeeline (Figure 1) that introduces a novel pipeline to enable users to write such micronotes and accurately capture the relevant context in the video to create the full (expanded) note. Micronotes being extremely brief, we need additional context to expand them properly (Monty, 1986). NoTeeline efficiently achieves this by utilizing the video transcript and previous examples of user writeup to decode the micronotes properly. Through a user study, we find that NoTeeline enables significantly faster and easier composition, and synthesis of notes than traditional notetaking tools. This work makes the following contributions:
  1. (1)
    We designed NoTeeline, a notetaking tool built around the concept of micronotes. NoTeeline enables users to write quick notes and then utilizes LLMs with relevant context to expand micronotes into full-fledged notes.
  2. (2)
    In-depth qualitative and quantitative evaluations that show the strengths and limitations of NoTeeline. We found that NoTeeline generates accurate notes with 93.2% factual correctness while significantly reducing the writing time up to 43.9%.
  3. (3)
    Based on our evaluation and user feedback, we provide design implication for designing better tools for effective Human-AI collaboration .
2. Related Works
2.1. Tools for Supporting Active Notetaking
Active notetaking has been long practiced as an efficient strategy to store and retain information (Peper and Mayer, 1978; Bauer and Koedinger, 2007). Past research has introduced several interactive notetaking tools to support live-notes writing (Weher and Poon, 1994; Hinckley et al., 2012, 2007; Liu et al., 2019; Tashman and Edwards, 2011; Tsai et al., 2024), inking (Forbus et al., 2011; Zheng et al., 2021; Cao et al., 2022), information scaffolding (Bernstein et al., 2008; Swearngin et al., 2021) and note organization (Spoerri, 2002; Meng et al., 2016; Subramonyam et al., 2020; He, 2014). NoteLook supports multimedia note taking in meetings in an integrated conference room (Chiu et al., 1999). VideoSticker proposes a visual notetaking tool for lecture videos utilizing object symbols to create multimodal, non-linear notes (Cao et al., 2022). AQuA introduces a question-answering interface that helps users to watch software tutorial videos to get quick information (Yang et al., 2024). Livenote allows students to take lecture notes and annotate lecture slides cooperatively (Kam et al., 2005). NoteLink allows students to link their hand-written notes with the relevant video so that they can replay it easily (Srinivasa et al., 2021). Although these tools help users to organize and link source materials faster, users still have to write the whole note manually, often requiring additional laborious revisions and manual effort, causing significant cognitive barrier (Piolat et al., 2005). On the contrary, NoTeeline focuses on automatically generating the text of full notes from micronotes to shift effort from text notation to capturing key ideas during active notetaking with users’ own note points. Instead of focusing on copying information word-for-word, users can concentrate on understanding and capturing the core ideas, minimizing the cognitive load associated with manual note-taking.
2.2. Supporting Shorthand Writing
Shorthand writing techniques have been popular with journalists for centuries (McCay, 2021), to support higher writing speeds without loss of information. Inspired by their success, many digital inking software drew inspiration from shorthand techniques (Leedham et al., 2009). This paved the way for a large body of research for gesture keyboard (Zhai and Kristensson, 2012). SHARK augments stylus keyboarding with shorthand gesturing by assigning a shorthand symbol to each individual word (Zhai and Kristensson, 2003; Kristensson and Zhai, 2004). Swipeboard encodes a sequence of gesture input into unique representations of each character to support faster typing in ultra-small devices such as smart watches (Chen et al., 2014). Both Swipeboard and SHARK report a higher writing speed compared to traditional keyboard. Padilla et al. introduce a deep learning-based methodology to automatically convert Gregg shorthand to English words (Padilla et al., 2020). Although these tools show the potential of shorthand-based writing techniques in the modern era, they specifically focus on gesturing interaction. Oftentimes, they also require users to learn existing shorthand symbols. NoTeeline provides users the flexibility to use their own abbreviated words and symbols without requiring users to learn new gestures to use our system. To be best of our knowledge, NoTeeline is the first tool to support micro-notetaking from videos and easy expansion of micronotes into full notes that are consistent with users’ corresponding writing styles.
2.3. Supporting Writing with LLMs
LLMs have emerged as popular tools for writing assistance because of their versatile generation capabilities. A wide body of research has focused on enhancement of LLMs as writing assistant through building novel interaction tools (Chen et al., 2019; Clark et al., 2018; Freiknecht and Effelsberg, 2020; Osone et al., 2021; Chung et al., 2022; Ghai and Mueller, 2021) and improving generation capacity of these models (Mysore et al., 2023; Li et al., 2023; Trajanovski et al., 2021; Tang et al., 2017a). Wordcraft allows users to write stories with LLM collaboratively with advanced features such as replacement or further prompting (Yuan et al., 2022). Ramblers allows users to create voice memos and collaboratively use LLM to edit and expand on the voice memos to create a gist (Lin et al., 2024). Pearl proposes a retrieval-augmented pipeline that utilizes previous user-authored documents to generate texts of specific user style (Mysore et al., 2023). ExpaNet retrieves relevant documents to an input query and expands it using the retrieved documents (Tang et al., 2017b). Outside academia, there is a growing number of commercially available tools for personal writing assistant as well (ProductHunt, 2024). While people frequently use these tools for assisting writing (Lee et al., 2022; Yuan et al., 2022), LLMs may struggle to preserve the style of experienced writers (Ippolito et al., 2022). Current research is increasingly directed towards personalized writing aids (Biermann et al., 2022). Compared to the above-mentioned tools, NoTeeline is designed for robust user customization of note generation from start to end via control of references (i.e: micronote) and writing style.
3. User Study Setup
Our study consists of two phases. First, we performed a needfinding study with 12 participants (§4) to understand the struggles they face in writing full notes in existing notetaking tools. Next, participants evaluated NoTeeline in one hour session. (§7). The entire session was recorded in Zoom. We recruited our participants through mailing lists and announcement boards. The participants are situated in North America. They received a \$30 Amazon gift card for their participation. The study was approved by the IRB committee in our home institution.
4. Needfinding
4.1. Procedure
We conducted a needfinding study to understand the current challenges users encounter with their current note-taking process. Specifically, we want to uncover underlying factors that make writing the full note difficult. We also want to understand whether users currently use AI-assistance for notetaking and explore the factors impacting their trust in AI generated notes. The results of this study informed the design choices we made for NoTeeline. Each session were approximately 20-30 minutes long.
4.2. Insights and Observations
Our participants reported using both manual note taking tools as well as AI assisted tools. The most common app used by participants was the built-in Macbook Notes app (n=9). Many participants use pen-and-paper notes and later digitize them manually. In terms of AI-based tools, some participants used ChatGPT (n=4) for summarizing meetings and video lectures. Others used Notion AI, Figjam AI, and Gemini plugin in Google Docs for specific usecases. Participants highlighted several challenges they face with their current note taking process. We summarize these findings into three key themes through iterative rounds of coding guided by qualitative data analysis (Yin, 2015):
C1. Users struggle to notate all desired information while taking notes from audio, video, and other real-time sources (×7)($\times$ 7)( × 7 ). Users mention resource constraints–time, attention, and mental load–in the practice of notetaking. P8 explains how their handwriting speed cannot match a professor’s lecturing pace. P7 details the additional difficulties imposed by being a non-native English speaker.
P11 finds it challenging to take notes while listening to podcasts, especially during physical activities like running on a treadmill. They describe the cumbersome process to jot down : ‘I have to slow down.. and then type it down in [a few words]… then I again need to speed up the treadmill and continue.’ The user emphasizes how the ad hoc nature of this process complicates effective note-taking, making it difficult to capture details in real-time.
C2. Lack of agency in AI-assisted notetaking tools (×$\times 6$). Users expressed the desire to maintain agency over AI assisted notetaking tools. This is important given concerns about LLM ‘hallucinating’ non-existent content (Ji et al., 2023). P1 says ‘I would, at the end of the day, want some kind of source of truth… based on what the notes are going to be’. They reported to prefer using automated note features ‘as a compliment to whatever that [they’re] using’ while ‘always going to take [their] notes no matter how’, because of the lack of trust in it. P2 further explains her thought - ‘I hope it’s a tool for me… to help me get better understanding of the video, not a tool for me to tell me what I should know, because I still want myself to have the control of what the tool is doing.’
C3. Complex features and workflow of current tools. (×7)($\times$ 7)( × 7 ). Participants highlighted that the complex features and fragmented workflows of current note-taking tools often hinder usability. One common challenge involved managing two separate applications—one for watching videos and another for taking notes—which made the process cumbersome. Many expressed a preference for integrating note-taking features directly within the video viewport, facilitating smoother interactions and easier linking of notes to specific video segments. This integration would also simplify the task of reviewing notes later on.
Furthermore, participants reported that the overwhelming nature of feature-rich tools like Notion discouraged full utilization. As P5 described, ‘I don’t end up using the extent of features that Notion has to offer… it’s so overwhelming .. so I think [I prefer] a simple interface but with the ability that as you go forward, you can start to explore more’. In other words, the features are not well-situated or contextualized within the workflow, and presenting them all at once can overwhelm users, preventing them from fully utilizing the benefits. In fact, it can deter users from engaging with the features altogether.
5. NoTeeline Interface
Figure 2. The NoTeeline user interface consists of (a) the Cues panel (outlined in blue) for review questions generated based on the notes, (b) the Notes panel (outlined in orange) for user notetaking and note expansion, and (c) the Summary panel (outlined in purple) for note summary. The layout is inspired by the Cornell method, as seen in the side-by-side comparison.
5.1. Design Goals
Based on our needfinding study, iterative prototyping and evaluation process, we propose three design goals for NoTeeline.
D1. Micronotetaking as the way to take notes from videos. The interface should enable the swift creation of micronote. Even if the micronote includes grammatical errors and/or is extremely short, the system should be able to comprehend it. While transforming micronotes, it should be aware of the context, i.e: the topic being discussed in the video. The system should enable iterative editing of micronotes. By focusing on micronotes, we want to tackle the difficulty in writing full notes users mentioned (C1).
D2. Maintain consistency with user’s style and preferences. The system should take into account the specific writing styles each user follows and remain consistent with their personal style. At the same time, the final note should exclude information users are not interested in. It should not include any hallucinated information, as desired by the participants in our findings (C2).
D3. Each component is linked to a specific note-taking procedure and shown accordingly. The system should link each component to a specific stage in the note-taking process—creation, editing, and revision. It should allow users to engage selectively with these features based on the current stage, providing ease of use throughout the note-taking journey (Pea, 1987; Nielsen, 2006) (C3).
5.2. Interface Components
We will now describe the components of NoTeeline built to support our design goals.
5.2.1. Onboarding Session: Collecting Personalized Examples.
We design an onboarding session to get reference examples of the user’s specific style (D2). Users watch three short clips (20-30 seconds each); and provide a keypoint as well as the corresponding full note for each clip (Figure 2). Once completed, the system will store these examples to use for note expansion afterwards. Users need to complete this session only once when they start using NoTeeline for the first time. This session also give users a first-hand experience of what a keypoint would look like and what kind of full note to expect from the system.
5.2.2. Notetaking Layout.
To construct the layout, we took inspiration from the Cornell method (Pauk and Owens, 2010), a popular technique for taking note (Saran et al., 2022; Piolat, 2007). We split NoTeeline layout into three regions – 1) Notes: the region for writing and organizing note; 2) Cues: questions or specific points that are helpful to review the note; and 3) Summary: main takeaway of the note in 5 sentences.
We also incorporated the video viewport on the top of the layout so that users can take notes while watching the video without needing to move between different desktop tabs (C3). It also progressively discloses the features as per users’ current need for note writing, editing or analyzing (D3). Figure 2 illustrates the notetaking layout and its different components.
Figure 3. Note expansion (some texts have been cropped for clarity).
5.2.3. Note Synthesis Operations.
We describe individual synthesis operations NoTeeline supports and their mechanism.
Micro Note Expansion. To construct the full note, we have an atomic expansion operation. For expansion, we utilize these pieces of information in the prompt: micronote, video transcript where the micronote was written, and reference examples from onboarding. The transcript helps to infer abbreviated symbols and fix potential typographical error (D1) (Trajanovski et al., 2021). Reference examples helps to understand users’ writing style (D2).
Note Organization by Theme. To rearrange the note points by emerging themes, we have an atomic order by theme operation (D1, D3). To order by theme, we pass the generated notes to the LLM and instruct it to return the output in a structured, parseable format by using the few-shot prompting technique (Jiang et al., 2023; OpenAI, 2023). We refer users to view the corresponding prompt Figure 11 in the Appendix to learn more about the structure we chose to use.
Figure 4. Note organization (some texts have been cropped for clarity).
Note Modification. To edit individual micronote or the expanded output from LLM and/or the automatically generated theme name (D3), users can click on the [Uncaptioned image] button next to each point.
Note Rearrangement. To enable users rearrange their notes (D3), we have a drag and drop operation. Users can drag individual note points and reposition them under a different theme using this operation. Users can also revert the note to its original serial by order by time operation.
Cue Question Generation. To enable users review their notes, we have a Cue Question Generation operation, which generates 5 questions on-the-fly. If users want new questions, they have the option to initiate Regeneration operation, which generates five new questions. The cue column is hidden by default to provide more space for writing to the users (D3). Cue question is generated in a similar structured fashion to order by theme operation. Please view the corresponding prompt (Figure 12) in the Appendix for details.
Summary Generation. To generate a personalized summary, we have a Summarize operation. The generated summary needs to reflect the user’s note as well as the video context (D2). Hence, we passed the full video transcript and expanded note to LLM for the summary.
5.3. System Walkthrough
We present a hypothetical scenario to illustrate how users might interact with NoTeeline. Kate is preparing a presentation on ‘Ocean Pollution.’ She uses NoTeeline to quickly go through a long 40-minute. She starts by completing the onboarding session and opens a new note named ‘Ocean Pollution Presentation.’ As she starts watching the video, Kate encounters a segment explaining the drastic increase in plastic pollution over the last decade. She quickly types ”plastic pol. $\rightarrow$” as a micronote. Once the video concludes, she presses the [Uncaptioned image] button, and NoTeeline elaborates her note into a detailed point, including the specific statistic: ”Plastic pollution in oceans has increased by 200% in the last 10 years.” This expansion allows Kate to easily integrate precise details into her presentation later.
Now, Kate needs to organize her information into themes so that she can design the slide decks accordingly. She clicks on [Uncaptioned image] and it organizes her notes into three themes - ‘Causes of Ocean Pollution’, ‘Effects on Marine Life’, and ‘Conservation Efforts’. While she is happy with the themes, Kate decides to further personalize the themes. She notices a point ’biodegradable alternatives to plastic’ that would be better suited under ‘Conservation Efforts’. Kate drags it under it and renames the theme to ‘Sustainable Solutions’ to better align with her presentation structure.
Feeling content with the organization of her notes, Kate clicks on [Uncaptioned image] to generate a personalized TL;DR summary for her conclusion slide. Finally, Kate wants to anticipate potential questions from her audience. She opens the [Uncaptioned image] panel. Attempting to answer from memory, Kate correctly addresses the first question but stumbles on the second, prompting her to revisit her notes for a quick recap. Satisfied with her understanding and the comprehensive coverage of her topic, she confidently moves on to finalizing her presentation, equipped with a well-structured and informative slide deck.
6. Prototype Implementation
NoTeeline is implemented using React.js and is hosted as a web app. All LLM-based operations and text manipulation are performed using OpenAI model, GPT-4-Turbo with hyperparameters set to temperature = 0.5, seed = 1. We chose GPT-4 over other LLMs due to its high-quality output and support for a longer size context window, especially suitable for longer videos (OpenAI, 2023). We use Zustand (Zustand, 2024) to manage the state of NoTeeline by storing user metadata locally.
7. User Evaluation
We invited 12 participants from our formative interviews to participate in a user study. Participants used both NoTeeline and a baseline to capture notes and evaluate their performance. Please note that participants were not exposed to NoTeeline’s system prior to the user study. They were divided into four subgroups to counterbalance and mitigate any ordering bias. Table 4 shows the distribution of participants across these subgroups. The test session was one hour long.
User Study Procedure. We first gave an overview of study procedures to the participants. Participants were given the NoTeeline web app link to browse from their laptops. They began with the onboarding session where we collected their personalized writing examples (§B.1). After onboarding, participants were given two videos (§B.2) to watch and take notes under the two study conditions. To prevent potential fatigue, participants were given a 5-minute break in between each video. After each video notetaking session, users were asked to complete a questionnaire (Hart and Staveland, 1988; Brooke et al., 1996; Sauro and Dumas, 2009). Finally, we asked participants try out NoTeeline with a video of their choice.
Baseline. To compare NoTeeline with a baseline, i.e., a basic notes app without any automated features, we used a simplified view of our interface by hiding all advanced features (i.e., the synthesis functions described in §5.2.3). This setting mimics the standalone macbook notes app, most of the participants reported to use. We also want to compare the utility between the personalized summary in NoTeeline and the automatically generated summary in existing tools. Hence, we kept the summary feature in the baseline which is similar in nature as NoteGPT (NoteGPT.io, 2024), i.e., a summary that is not conditioned on user notes.
8. Results and Findings
Micronote Expanded note
RNNs are unrolled l to r or opp Recurrent neural networks (RNNs) can be unrolled either from left to right or from right to left, encoding linear locality.
find the average / adding vectors together Average of vectors -> add vectors together -> find mean value.
red -> all similar reds Red color categorization includes various shades that are perceived as similar.
Cites difficulties for BVI people
The speaker cites the difficulties faced by blind and visually impaired individuals, such as the inability to independently
read books, magazines, or use computers and smartphones without assistance.
sauce: gehra?
The sauce, known as garum, was a staple in Ancient Greek and Roman cuisine, valued for its strong umami flavor and
produced in large quantities throughout the Mediterranean.
both rich and umami Tomatoes, like fish sauce, are rich in glutamate which gives them a strong umami flavor.
css property –> native styling Custom properties in CSS can be used to create native styling options for web design.
met The Metropolitan Museum of Art is located on Fifth Avenue and spans five blocks.
))) - russian In Russian, multiple closing brackets ”)))” indicate varying degrees of happiness in a smiley face.
pipe –> high to low pressure Water in pipes flows from areas of high pressure to low pressure, similar to how electrons move from high to low voltage.
equipment durability Cycling pumps on and off to meet variable demand increases wear and tear, affecting equipment durability.
slow - reducing fat mass. fast - reducing muscule mass Slow dieting reduces fat mass, while fast dieting leads to muscle mass loss.
1 enery deficit Sam creates an energy deficit by consuming fewer calories and increasing exercise, leading his body to break down glucose stores.
loose fast? < find healthier life style Losing weight quickly isn’t ideal; instead, focus on discovering a lifestyle that promotes personal health and happiness.
Applied Changes: X    Abbreviation Expansion X      X    Detailed Information Extraction X      X    Personal Writing Style Imitation X      X    Personalized Symbolic Interpretation    X      X    Typo Fix
Table 1. Example micronotes written by our participants and the corresponding expanded notes in NoTeeline. The relevant portions in the micronote and expanded note are highlighted for readability. The final notes can properly expand the abbreviation as well as symbols, extract detailed information, mimic personal writing style, and fix typographical errors.
Figure 5. Distribution of pause and skips used by each participant in NoTeeline (left) and Baseline (right). In NoTeeline, the count of pause and skip is significantly less than the baseline. An empty bar depicts that the user did not pause or skip at all.
We describe the results of our user evaluation, both quantitatively and qualitative showing the effectiveness of NoTeeline in notetaking.
8.1. NoTeeline helps Users to Write Notes Efficiently with Reduced Disruption
Users efficiently captured the notes with less text in NoTeeline. We report the average length of notes written by users in NoTeeline ( $\mu=32.11,\sigma=12.33$), shorter than that in baseline ($\mu=60.64,\sigma=16.94$) (Figure 6(a)).
Users took more notes in NoTeeline. The total number of notes is more in NoTeeline ($\mu=18.75,\sigma=5.69$) compared to baseline ($\mu=14.75,\sigma=2.59$) (Figure 6(b)).
Users took a shorter time to write notes in NoTeeline. Compared to baseline ($\mu=27.79,\sigma=8.25$), the notetaking time was 43.9% shorter in NoTeeline ($\mu=15.59,\sigma=5.4$) (Figure 6(c)).
Users faces fewer disruption in watching the video in NoTeeline. Users paused or rewound less with NoTeeline ($\mu=0.56,\sigma=1.86$) compared to baseline ($\mu=4.94,\sigma=5.97$) when they were writing notes (Figure 5).
The results indicate NoTeeline enabled users to capture notes more efficiently by reducing the amount of text they need to write to express their thoughts. This also helps users to deeply concrete without disrupting their focus by pausing or rewinding. P4 directly speaks to this advantage: ‘I felt like I had to.. take every note [in Baseline].. you saw me like pause and rewind… Whereas in [NoTeeline ], I felt like I could just watch it all the way through and not have to rewind.’ The higher frequency of notes also show that NoTeeline does not hinder the notetaking process in any way, rather it helps them to write it with less effort. Similar to how LLM writing tools facilitate more refined thinking (Nakazawa et al., 2022), NoTeeline enables users to engage more actively with video content by allowing them to write notes with less effort.
(a) Average length of note point.
(b) Total count of note point.
(c) Average time per note point.
Figure 6. Comparison of (a) note length, (b) count, and (c) writing time for each participant between NoTeeline and baseline. Majority of the users took higher amount of notes in NoTeeline in less time.
Figure 7. User study results (N=12). Boxplots from left to right: (a) workload measured with NASA TLX (Hart and Staveland, 1988) (7-point Likert scale, lower is better), (b) usability measured with SUS (Brooke et al., 1996) (7-point Likert scale, higher is better), (c) task completion time in seconds (lower is better).
8.2. NoTeeline supports a Wide Range of Notetaking Pattern and Associated Transition
Table 1 presents some example micronotes our participants generated and their corresponding expanded notes generated by NoTeeline.
Patterns: Participants utilized a variety of patterns when constructing their micronotes. Symbol usage was frequent, with ‘-¿’ commonly used to indicate a process. ‘?’ was often employed to denote uncertainty or to highlight areas where participants sought further clarification. Abbreviations were widely used as well. For instance, P1 used the acronym ’BVI’ to represent the ’blind or visually impaired community.’ P2 employed the numbers ’1’ and ’2’ to refer to characters ’Sam’ and ’Felix’ from a video. Some participants shortened words by using the first letter or a few subsequent letters. For example, P8 wrote ’l to r’ to represent ’left to right.’
Transitions. NoTeeline supports five types of transitions during note expansions. 1) Abbreviation Expansion: Converts abbreviations into their full form; 2) Detailed Information Extraction: extract missing information from transcript required for note expansion. 3) Personal Writing Style Imitation: mimic the writing pattern users usually use in their writeup; 4) Symbolic Interpretation: accurately decodes the symbols in user-specific writing style and integrates them into the expanded notes; 5) Typographical Error fix: corrects spelling mistakes or missing letters in the text.
8.3. Comparison of NoTeeline and Baseline
8.3.1. Workload.
Using the Shapiro-Wilk test of normality, the differences between baseline and experimental condition responses for three of the four NASA-TLX questions (Hart and Staveland, 1988) on workload significantly deviate from a Normal distribution. Hence, we compare the differences in workload through a Wilcoxon One-tailed Signed-Rank Test (Neuhäuser, 2011) with a Bonferroni-corrected (Weisstein, 2004) significance level of 0.0125 per condition. Participants report significantly lower results for Mental Demand, Effort, and Frustration ($p=0.0123,p=0.0024,p=0.0047$). We did not find statistically significance for higher Performance (p = 0.0803), this may be attributed to users’ general familiarity with notetaking, self-assessments of performance were already high in the baseline condition ($\mu=5.58$ in the baseline versus a slightly higher average of $\mu=6.25$ in the experimental condition on the 7-point Likert scale) and results may be subject to the ceiling effect (Hancock, 1989) (5 out of 12 participants gave the maximum 7 rating).
8.3.2. Usability.
Usability was measured via an adapted version of SUS (Brooke et al., 1996)) and yielded an overall average of 80.972, which corresponds to an A grade according to the Sauro–Lewis curved grading scale (Sauro and Lewis, 2016). We note that the average baseline SUS values are also high at 80.833. Thus, our experimental features do not seem to detract from the inherent usability of notetaking methods more familiar to users. Further, our baseline condition already provides salient advancements to basic notetaking by combining the video source and notetaking area with other AI-assisted features—all of which received generally positive comments from participants—into one area. This is salient as P2, P3, P4, P10, P12 brought up issues of screen space economy and the importance of a proper split screen setup during notetaking (e.g. P12 mentions advantages of our setup over typical split screen set-ups for notetaking, which pose difficulties for tab navigation).
8.3.3. Task Completion Time.
The differences in task completion times per participant across conditions pass the Shapiro-Wilk test of normality ($W=0.8665,p=0.0590$). Hence, we compare the differences in task completion times (seconds) through a parametric paired t-test. Participants report a significantly lower task completion time when using NoTeeline ($\mu=207.7,\sigma=17.25$) compared to the baseline ($\mu=322.36,\sigma=103.27$) ($t(11)=3.6,p=0.004$) (Figure 7(c)).
Figure 8. Participants’ responses to utility and usability of NoTeeline interface and various features (7-point Likert scale, higher is better).
8.3.4. Feedback on NoTeeline features.
In figure 8, we show participants rating for individual features in NoTeeline ($\mu=5.8,\sigma=1.15$) (7-point Likert scale, higher is better). Overall, NoTeeline was received positively by the users. Please see §8.7 for more detailed feedback on NoTeeline’s features.
8.4. Note quality of NoTeeline
Linguistic Property. We report the linguistic quality (in terms of grammaticality, non-redundancy, focus) of notes using GRUEN (Zhu and Bhat, 2020). The result (Table 5 in Appendix) indicate that the points written using NoTeeline ($\mu=0.84,\sigma=0.07$) are of higher quality (i.e: less grammatical less, less redundant and more focused) than the baseline ( $\mu=0.56,\sigma=0.17$) ($p<0.0001,t(11)=10.87$).
Correctness. To ensure our automatically expanded notes are indeed correctly extracting the information from the transcript and not hallucinating any misinformation (C2), we compute the factual correctness of the generated note using HHEM (Vectara, 2024; Zhang et al., 2023; Azamfirei et al., 2023). We noticed a Hallucination Rate = $6.8\%$ and a Factual Consistency Rate = $93.2\%$. This indicates that our generated notes are correct and consistent $93.2\%$ of the time.
User evaluation. We also collect the self-reported scores by the users. On average, users rated the expanded note, $\mu=6$ ($\sigma=1.044$) on a scale of 1 to 7 (higher is better) (Figure 8).
8.5. Consistency of Micronote and its Corresponding Expanded Note
To ensure our expanded notes are consistent with the main content of the micronotes, we measure the widely used sentence representation similarity, SBERT (Reimers and Gurevych, 2019; Ethayarajh, 2019; Gao et al., 2019). We find that they have a 58% similarity (avg. score = 0.58). We also examine the similarity of expanded note and the transcript and found it to be 42% (avg. score = 0.42). This results show that the expanded notes are more similar to micronotes than the transcript, indicating that our expanded notes prioritize the micronotes’ contents and do not merely replicate the video transcripts, which will be undesirable.
8.6. Consistency of Summary and Note
To measure if the generated summary properly summarizes on the points mentioned in the note, we measure sentence representation similarity (SBERT) between the note and generated summary(Reimers and Gurevych, 2019; Ethayarajh, 2019; Gao et al., 2019). Compared to the baseline (i.e: NoteGPT (NoteGPT.io, 2024)), NoTeeline summary is 11% more relevant to the user note (Table 2). Table 6 shows few examples of summary and expanded note pairs in the Appendix.
P1 P2 P3 P4 P5 P6 P7 P8 P9 P10 P11 P12 avg.
Consistency of Micronote, transcript & corresponding expanded note (§8.5)
Micronote &
Expanded Note
0.6 0.41 0.73 0.75 0.63 0.58 0.55 0.6 0.48 0.43 0.57 0.63 0.58
Transcript &
Expanded Note
0.47 0.41 0.42 0.42 0.42 0.39 0.42 0.42 0.44 0.46 0.39 0.41 0.42
Consistency of Summary & Note in two conditions (§8.6)
Baseline 0.81 0.53 0.81 0.61 0.64 0.76 0.33 0.67 0.65 0.71 0.44 0.72 0.64
NoTeeline 0.87 0.51 0.9 0.63 0.91 0.77 0.67 0.8 0.68 0.7 0.76 0.82 0.75
Table 2. Technical Evaluation Result (measured as sentence representation similarity (Reimers and Gurevych, 2019), higher is better)
8.7. How NoTeeline Features Assist in Notetaking
8.7.1. Impact of Note Expansion.
All 12 participants reviewed some aspect of the expanded note favorably. All critical feedback, present in four of the responses, focused on the potential for verbosity and redundancy in the expanded notes (refer to §8.4 for demonstration of how NoTeeline still improves upon the baseline for measures of redundancy). Still, nine reviews expressed appreciation for the quality/level of detail in the expanded note, and five reviews brought up that the expanded note captured details/context they would have otherwise missed, due to such reasons as cognitive and time constraints (e.g. P9: focusing on the key words is less mentally draining and would be helpful in keeping my attention span for a longer period of time’; P6: ‘there’s like a lot of things that I heard I probably wouldn’t have to have time to type it all out, but it was able to expand on everything’). Eight participants also specifically remarked that the expanded note reflected their writing style. P11 says ‘If my note was seven words or eight words then it went like one and a half lines if mine was small and then it was also like not trying to bring in too much content so that I felt it was trying to match with my style of what I’m trying to put’.
8.7.2. Impact of Generated Themes.
11 out of 12 participants reviewed some aspect of the generated themes positively. P8 believes the impact of Order by Theme would become more evident in longer videos. P10 discusses how it would have greater impact in a video with more discursive content. In that regard, P11 expresses excitement about using the Order by Theme in such a context ‘where the conversation may not usually flow through on one topic’. Five reviews specifically point out the utility of the Order by Theme feature for aiding review, retrieval, and organization of content.
8.7.3. Impact of Cue Questions.
11 out of 12 participants reviewed some aspect of the cue questions positively, with critical feedback from five participants tending to focus on increased cue question complexity. P9 identified that multiple choice question may not be applicable for ‘concepts [that] are more open to interpretation’. Six participants specifically brought up positive impressions of the cue questions aiding in the video/note review process, four participants mentioned that they viewed the cue questions as a way to identify potential shortcomings in their notes, and three participants highlighted perceived relevance of cue questions to the video content.
8.7.4. Impact of Summary.
11 out of 12 participants reviewed some aspect of the summary feature positively. Four responses brought up satisfaction with the summary’s ability to capture the gist/important contexts from their notes. Three responses mentioned a preference for notes versus the generated summary for reviewing purposes: P1 prefers reviewing ‘small details’ while P3 and P6 find the length too long. However, opinion on length is not uniform, as five responses (P5, P7, P8, P10, P12) reflected positively on the length of the summary. Four responses described a desire for more formatting features, such as bolded or highlighted text.
8.7.5. General Reception of AI Assistance
We conclude thematic analysis with a general review of reception to NoTeeline’s AI-assistance during notetaking. Across the user study, ten participants specifically addressed the relationship between their labor and the AI features in NoTeeline impacting their workflows. P1 speaks to the ‘division’ of labor in which they capture the ‘smaller details’ while the ‘machine… take[s] over the higher level summary.’ P1, P3-P4, P6-P8, P10-P12 speak to how the automated features distributed throughout NoTeeline provide enhancement or safeguard beyond their individual labor–whether capturing missed details/context, correcting typographic and grammatical errors or otherwise refining their writing, or surfacing unrecognized themes or shortcomings in their understanding of the material.
8.8. NoTeeline Enables Efficient Note-taking across Diverse Domains
During the freeform test (§7), Participants reported and used NoTeeline in a wide range of video categories. All 12 participants wished to continue using NoTeeline in future. P7 continued using NoTeeline for annotating interviews collected as part of their research even after our study session. P9 mentions the desire for integration with a current Learning Management System, Canvas, used at their university. Most participants thought NoTeeline would be most useful for long-form videos where there is a lot of information to capture rather than short videos which may not that essential to take notes on. The top categories they reported are: 1. Educational Video Content (×$\times 8$) 2. Professional Podcast and meetings (×$\times 3$) 3. Thematic analysis of research videos (×$\times 3$) 4. Personal Learning and Self-Education from how-to tutorials (×$\times 2$) 5. Financial and Investment Learning (×$\times 1$).
Even when features were not immediately relevant to their current needs, participants reported potential uses for others or themselves at other time points in their lives. P12 says the cue questions would be useful for ‘my (their) friends at med school’ where ‘the details get really granular and… the names of things can can make or break like the answer to a question.’ P6 mentions how, even though they are no longer a student, they would have ‘definitely want[ed] to use [NoTeeline ] with [their] lectures… freshman and sophomore year when classes were online.’
8.9. Error Analysis
Error in LLM generation. GPT-4 sometimes fails to generate the requested note. For example, P7 wrote a micronote for noting down her thoughts that is not directly discussed in the video. The model replied with - ‘Please provide the transcript related to the keypoint so I can assist you in creating a note.’. The occurrence of such error is 0.72% (4 out of 559 total calls).
9. Discussion
9.1. Implication of NoTeeline in Modern real-time Notetaking
Building upon the established concept of micronotes, NoTeeline supports real-time notetaking from videos. By leveraging large language models (LLMs), NoTeeline reduces the manual effort required for writing full notes, allowing users to focus on capturing core ideas rather than transcribing content verbose.
Our study shows that users were able to produce an equivalent or greater amount of information in their notes using fewer words and less time. This suggests that micronotes, when combined with NoTeeline’s automation, remain a highly effective method for real-time notetaking without requiring extensive user involvement or high precision. The system helps users concentrate on the material being discussed rather than the act of writing, addressing common pain points associated with the overwhelming complexity of existing note-taking tools.
While concerns may arise about potential disengagement with the material due to automation, we argue that NoTeeline enhances rather than detracts from the cognitive process. Users still initiate and guide the note-taking process, with NoTeeline simply assisting in transforming their initial input into structured notes. This elucidates the workflow as well as promotes deeper engagement by freeing users from the mechanical task of writing. Prior research on LLM-based writing assistants also report that automating the writing experience does not necessarily hinder the main objective but rather enables more refined thinking (Nakazawa et al., 2022; Göldi et al., 2024). Furthermore, NoTeeline incorporates a cue question feature inspired by the Cornell method, which has been shown in previous studies to improve long-term knowledge retention (Kornell, 2014; Evans and Shively, 2019; Yuniarti and Trisnawati, 2018).
9.2. Design Implication for Human-AI Collaborative Tools
In human-AI collaboration, trust plays a critical role in determining how effectively users engage with AI systems. In our Needfinding study, we likewise uncover user concerns about LLM hallucination and control in the domain of note-taking (e.g. P1’s and P2’s quotes in theme C2, §4.2). Research shows that trust in human-machine teaming is built through the reliability, transparency, and explainability of AI systems (Alhaji et al., 2020). By controlling the precise moments of note-taking instead of relegating the summarizing task entirely to note-taking software, we offer greater transparency and specificity in the note-taking process. This is further supported by the low hallucination rates we uncovered in Section 8.5. This suggests that ensuring users can trust the AI to expand micronotes in a way that aligns with their personal keypoints and style is essential for fostering productive collaboration. Additionally, the ability to modify or intervene in the AI’s output has also been shown to increase trust (Jeung and Huang, 2023). In our tool, we achieve this via our note editing feature and the involved nature of the general micronote expansion process.
Further, we uncover the benefits of designing for implicit guidance by aligning the processes of task completion and training processes. In our system, note-taking serves a dual purpose — simultaneously achieving the note-taking task and providing continuous guidance to the LLM to improve user experience in a positive feedback loop, reducing needing additional external signals.
10. Limitations and Future Work
Through NoTeeline’s testing phase, we uncovered potential areas for future research. First, we could further organize notes in multi-level, hierarchical list formats to improve the note review experience. Second, we can incorporate support for rich text formatting (i.e: text highlighting, support for different font styles and sizes), which can be implemented trivially by incorporating rich text formatting package (quilljs.com, 2019). Third, we can potentially reduce the response time of GPT-4 by parallel computing or pre-computing future operations. Some recent works streamline the model output to reduce the wait time (Lin et al., 2024), but we did not prefer it in NoTeeline. Our concern was that any inaccuracies in the streamlined output might complicate error correction, as the information is already presented to users, potentially compromising the integrity of the user experience. We can also consider the inclusion of non-textual note elements, such as images and drawings. Finally, as surfaced by participants multiple times throughout the user study, a promising direction can be evaluating the potential learning outcomes of NoTeeline, i.e: to what extent it can enhance academic performance and learning retention. Since we did not specifically consider academic contexts in NoTeeline, our current focus has been on its efficacy for swift, personalized notetaking (Lin et al., 2004). More generally, future work may further explore the complexities of AI-assisted notetaking in contrast to other writing domains (Bedington et al., 2024).
11. Conclusion
Active notetaking from videos remains challenging. While existing tools can automatically generate notes, they can not capture the stylistic nuances users care about. Our needfinding study reveals that users often do not prefer an automatically generated summary due to concerns about accuracy and lack of customization. On the other hand, NoTeeline leverages rich contextual information to generate more accurate and more personalized notes. Our user study reveals that NoTeeline helps write notes faster with $93.2\%$ correctness and achieves closer stylistic characteristics to users’ handwritten notes. We demonstrate the utility of NoTeeline in a variety of usage scenarios users reported and provide design implications for future works in human-AI collaboration.
Societal Impact
By helping users take notes faster, our tool has potential to make note-taking more accessible for individuals who might otherwise struggle due to reasons such as language fluency (e.g. P7’s comments about difficulties of being a non-native English speaker). While there exist general arguments about how over-reliance on AI-generated content may lead to passive engagement or diminish personal organizational skills, we show how our tool exists to amplify not override user engagement by generating content driven by direct user input. We expect this should also decrease the risk of AI biases that might occur in more fully generated content (e.g. information prioritized during note generation). Finally, while the mechanisms of our tool are not necessarily in themselves geared towards learning outcomes, facilitating faster and more extensive information capture–in a personalized way via a compact UI–has the possibility to help users organize, synthesize, and engage with content in ad hoc ways towards educational, professional, and personal use cases.
Acknowledgements
We thank Dan Saffer and Chris Maury for their valuable feedback on the design of NoTeeline, which contributed significantly to its development. We also appreciate the insightful comments and suggestions provided by Pranav Khapde, Nazmus Saquib, and Yi-Hao Peng on earlier drafts of this work.
Appendix A Participants’ Information
Table 3 shows the demography and occupation of our participants. Table 4 reports the subgroup division of the participants into four groups. Figure 9 shows the reported tool used by our participants as a tree map.
ID Gender Occupation
P1 Male Graduate Student
P2 Female Undergraduate
P3 Female UX Designer
P4 Female Undergraduate
P5 Female Masters Student
P6 Female Student
P7 Female Graduate student
P8 Female Masters Student
P9 Male Undergraduate
P10 Female Student
P11 Male Product Manager
P12 Male Software Engineer
Table 3. Demography of user study participants.
Group No. # of Participants Tool_1 Video_1 Tool_2 Video_2
1 3 (P1, P5, P9) NoTeeline url1 Baseline url2
2 3 (P2, P6, P10) Baseline url1 NoTeeline url2
3 3 (P3, P7, P11) NoTeeline url2 Baseline url1
4 3 (P4, P8, P12) Baseline url2 NoTeeline url1
Table 4. Distribution of participants across four subgroups.
Figure 9. Different tools and techniques participants reported to use for note-taking in their daily lives.
P1 P2 P3 P4 P5 P6 P7 P8 P9 P10 P11 P12
Baseline $0.69\pm 0.18$ $0.56\pm 0.15$ $0.55\pm 0.17$ $0.62\pm 0.24$ $0.47\pm 0.2$ $0.60\pm 0.15$ $0.5\pm 0.14$ $0.4\pm 0.18$ $0.57\pm 0.16$ $0.61\pm 0.15$ $0.67\pm 0.18$ $0.53\pm 0.18$
NoTeeline $0.86\pm 0.05$ $0.84\pm 0.04$ $0.84\pm 0.04$ $0.83\pm 0.07$ $0.86\pm 0.07$ $0.83\pm 0.06$ $0.73\pm 0.19$ $0.86\pm 0.05$ $0.83\pm 0.11$ $0.84\pm 0.06$ $0.85\pm 0.04$ $0.86\pm 0.04$
Table 5. Gruen scores of each participant
Appendix B Video Selection Criteria
B.1. Video Clip Selection for Onboarding Session
We carefully selected three videos for the onboarding from three different domains: Science, Psychology, and blogs. We extracted 15-20 seconds snippet from each video and stored the timestamps. The videos are listed below: 1) How Mitochondria Produce Energy (Science); 2) How to get motivated even when you don’t feel like it (Psychology); 3) Apple Vision Pro VS Meta Quest 3! HONEST Comparison! (Blog).
B.2. Video Selection Criteria for User Study
We picked two videos for the two conditions. The videos met the following criteria: 1) not too long ($<5$ minutes); 2) information heavy where users are exposed to scientific concepts (– so that they feel the need of taking note); 3) interesting to a wide range of audience. Based on these criteria, we selected two TED-Ed videos entitled ‘How memories form and how we lose them’ (url1) and ‘Is it possible to lose weight fast?’ (url2). Both videos relate to Human Biology, are ~4 minutes long, and have been viewed over 2 million times, reflecting their wide appeal.
If time permitted, we also encouraged participants to try out NoTeeline with a video of their choice to stir personal interest. For this, we explicitly asked participants to freely explore NoTeeline and observed no carry-over biases. Similar ordering of reproduction task and free creation task has been adopted by previous HCI works (Xia et al., 2018, 2016; Cao et al., 2023).
Expanded Note Summary
Encoding is done with bidirectional LSTM, and the attention mechanism allows
the decoder to treat encoded representations as a form of memory to reference
and highlight what’s important at any time.
In the lecture, the use of bidirectional LSTM in encoding is highlighted, with the attention mechanism enabling the decoder to utilize encoded representations as a memory form, focusing on relevant parts at any given time. This approach is foundational in sequence to sequence models, particularly in creating end-to-end differentiable machine translation systems, where an encoder and decoder framework is pivotal. The lecture also touches upon the nature of recurrent neural networks (RNNs), which can be unrolled in both directions, thus capturing linear locality and maintaining a computational complexity that is linear with respect to sequence length. This linear encoding of locality by RNNs contrasts with the quadratic computational complexity challenges faced by transformers, underscoring the evolution and adaptation within NLP technologies.
Sequence to sequence models employ an encoder and decoder framework
to handle tasks like building an end-to-end differentiable machine translation
system.
Recurrent neural networks (RNNs) can be unrolled either from left to right or
from right to left, encoding linear locality.
Recurrent neural networks encode linear locality with a computational
complexity that is linear in terms of the sequence length.
Table 6. Expanded notes and summary generated by GPT-4 using these notes. The relevant portions in the expanded notes and summary have been highlighted for readability.
Appendix C Relevance of expanded note and summary
Table 6 shows how the summary is conditioned on user notes and includes the relevant portions from the notes.
Appendix D Prompts used for NoTeeline
D.1. Prompt for Note Expansion
Figure 10 shows the prompt that we use for expanding the micronotes of the user. It specifies some rules explaining what kind of expanded notes are expected. The rules ask to resolve any grammatical mistakes, generate a single sentence as the expanded note point, and avoid some specific phrases to start the expanded sentence. The onboarding transcript, onboarding micronotes, and the onboarding full notes are supplied as examples to the prompt. Moreover, the transcript of the YouTube video and the microntes of the user are also passed to the prompt for expansion.
D.2. Prompt for Note Organization by Theme
This prompt can be seen in Figure 11. It explains the structure of the theme that is expected as a response. As an example, some points and their corresponding themes are incorporated into the prompt for better theme generation. Like the note expansion prompt, some rules have also been mentioned here. These rules tell how many points are to be grouped into one theme topic and that all points should be assigned a detailed theme topic.
D.3. Prompt for Cue Question
The prompt here in Figure 12 asks to generate five multiple-choice questions based on the expanded note that has been provided. It also shows in what format the response is expected. A topic and five multiple-choice questions have been added to the prompt.
D.4. Prompt for Summary
The prompt for summary in Figure 13 consists of a context and some keypoints. The context here is the preprocessed transcript in Langchain and the keypoints are the expanded note points. It strictly asks to generate a summary in 4 sentences and use the context only for additional information.
Appendix E Questionnaire for NeedFinding Study
We asked the participants the following five questions in the Needfinding study:
  • What is your typical notetaking process when watching online informational videos (e.g., podcasts, lecture, tutorials)?
  • What are your main pain points with the current notetaking process?
  • Do you use any existing AI software to assist with your notetaking? If so, please briefly describe their pros and cons.
  • Are there any additional note taking features that you would like to have in the tools you are using now?
  • If there is an AI assisted system that could take your manual notes and organize them automatically for you, would you like to use them?
{mdframed}
[] I want you to act as a personalized note-taking assistant. Users will give you a keypoint and the youtube transcript. Your task is to expand the keypoint into a note point, by taking additional context from the transcript. The note should be a full sentence in simple English. Follow these rules:
1. Resolve any typos or grammatical mistakes that arise in the keypoint.
2. The note should not be longer than 1 sentence.
3. Remember that the keypoint can be very abstract and as short as an abbreviation. Use the transcript to get additional information to ensure a good quality note expansion.
4. Just write a single note point, users will request repeatedly for new points they want to add.
5. Write it in a way a user would write in a notepad. Do not use sentences such as ’This video talks about...’, ’The speaker explains..’ etc.
Make sure that the note aligns with the user’s writing style, so that they can read it easily. Use the same writing style as shown below. Here are three examples:
Transcript: Onboarding Transcript
Keypoint: Onboarding Micronotes
Note: Onboarding Full Note
The keypoint refers to the high-level keypoint provided by the user and your task is to write a full ’Note’ point. Make sure that your expanded note point matches the writing style of ’Note’ in the provided examples. Expand the provided keypoint into a one sentence note.
Transcript: Input Transcript
Keypoint: Input Micronotes
Note:
Figure 10. Prompt used for note expansion
{mdframed}
[] Given a list of points, Your task is to perform topic modeling over them. Arrange the points into topics and provide a name to each individual topic. Please mark the topic within <Topic></Topic>tags. the points are marked with <p>tag as well. Follow these rules:
1. Each topic should have more than one points.
2. Every point should be assigned under one topic. No point should be unassigned.
3. The topic should not be very high-level, make them as detailed as possible.
Here is an example:
Input points:
1. New York City is home to a vibrant mix of cultures, with communities from nearly every corner of the globe, offering an incredible range of languages, cuisines, and cultural practices. .. (cont.)..
Answer:
<Topic name="Cultural Celebration in New York City">
<p>New York City is home to a vibrant mix of cultures, with communities from nearly every corner of the globe, offering
an incredible range of languages, cuisines, and cultural practices.</p>.. (cont.)..
</Topic>
Input points: Expanded Points
Answer:
Figure 11. Prompt for Note Organization by Theme
{mdframed}
[] System prompt:
Given a topic description, your task is to generate five multichoice question with answer. Please mark the question within <Question></Question>tags, individual choices within <Choice></Choice/>tags and answer within <Answer></Answer>tags. Make sure not to put the write choice in the same choice option, randomly assign it within A, B, C or D. Here is an example:
Topic: Resilience refers to how well you can deal with and bounce back from the difficulties of life. It can mean the difference between handling pressure losing your cool. Resilient people tend to maintain a more positive outlook and cope with stress .. (cont.)..
<Question>What does resilience refer to?</Question>
<Choice>A. Dealing with difficulties by losing your cool</Choice>
<Choice>B. Bouncing back from the challenges of life</Choice>
<Choice>C. Avoiding stressful situations altogether</Choice>
<Choice>D. Ignoring problems and hoping they go away</Choice>
<Answer>B. Bouncing back from the challenges of life</Answer>
User prompt:
Topic: Expanded Note
Additional Context to get extra information for the topic: Preprocessed Transcript in Langchain
Figure 12. Prompt for Cue Question
{mdframed}
[] I will give you a context and some keypoints, Your task is to summarize the keypoints in 4 sentences. Focus on the keypoint, only use context if you need extra information:
Context: Preprocessed Transcript in Langchain
Keypoints: Expanded Points
remember not to make it too long. Do not mark the sentences with 1, 2, etc.
Figure 13. Prompt for Summary
P1 P2 P3 P4 P5 P6 P7 P8 P9 P10 P11 P12
Hallucination Rate 4.76 5.88 11.11 6.67 6.67 9.09 0.0 0.0 10.71 13.33 10.0 3.33
Factual Consistency Rate 95.24 94.12 88.89 93.33 93.33 90.91 100.0 100.0 89.29 86.67 90.0 96.67
Table 7. Correctness of generated note for each participant in terms on Hallucination Rate (%, lower is better) and Factual Consistency Rate (%, higher is better)